10.g of a certain metal absorb 40. cal of heat and the temperature is abserved to rise from 20 *C .The specific heat of this metal is _ cal/g *C

100.cal of heat are addad to 10 g of ethanol( 0.581 cal/g *C) origginally at * C. The final temprature is_

How many joules of heat are reguired to raise the temperature of 35 ml of water from 25 *C to 100 *C?

still waiting for anwers
Answer #1

m = 10g

   q = 40cal

T   = 200C

q = mcT

c   = q/m*T

        = 40/(10*20)   = 0.2cal/g-0C

2. t1 = 15^0C

    m = 10g

    c   = 0.581cal/g-c

q = 100cal

q   = mcT

q   = mc(2t-t1)

100= 10*0.581*(t-15)

t = 32^0C

3.

mass of water = volume * desnity

                       = 35*1 = 35g

no of moles of water = 35/18    = 1.94moles

q1   = mcT

        = 35*4.184*(100-25)

        = 10983J

q2   = nH vap

       = 1.94*40.79

        = 79.1326Kj    = 79132.6J

total heat q = q1 + q2

                     = 10983 +79132.6

                      = 90115.6J